OpenStudy (anonymous):

Solve and check 6+5/4 x=-4 please show work

OpenStudy (anonymous):

6+5/4x= -4 -6 -6 5/4x= -10 5/4 5/4 x= -8

OpenStudy (anonymous):

well like i subtracted 6 from both sides. then i divided each side by 5/4. and i got -8

jhonyy9 (jhonyy9):

6 +5/4 x = -4 5 --- (x) = -4-6 ------- 5x= -40 --- x= -40/5 --- x= -8 4